# Fault-tolerant hyperbolic Floquet quantum error correcting codes

### v3

@@ 3 Error Models and Decoding Method
We employ two error models. A phenomenological model comprising unbiased depolarization errors affecting individual physical qubits and measurement errors associated with the weight-two measurements of the checks, and an entangling measurements model which imitates circuit-level noise for platforms with native two-qubit measurements, such as Majorana fermions. Considering other noise models and optimizing the code/decoder to the noise is an interesting avenue for future work.
@@
Inferred values of the plaquette operators, called plaquette syndromes or simply syndromes, construct a ${(2+1)D}$ space-time lattice that we refer to as the syndrome lattice. In the absence of errors, consecutive inferences of each plaquette syndrome yield the same value. Conversely, an updated plaquette syndrome that differs from its previous value implies the presence of an error (or errors) affecting that syndrome. We call these events detection events, and they are used in identifying the occurred errors.
Each single-qubit Pauli error, depending on the type of Pauli operator and the last round of measurements, yields either two or four detection events. We call Pauli error with two detection events simple Pauli errors and those with four detection events compound Pauli errors. Two-qubit Pauli errors give rise to the detection events corresponding to their constituent single-qubit Paulis, and measurement errors are associated with four detection events. (see Fig. 8 and Fig. 9.)
@@ 4 Fault-Tolerant Memory Threshold
For each hyperbolic Floquet code listed in Table 1, this analysis involves sampling instances of decoding scenarios. In each decoding scenario, depolarization and measurements error channels with strength $p$ (as defined in Sec. 3) are applied during $3d$ rounds of check measurements, corresponding to $d$ Floquet periods, where $d$ is the code distance. Repetition of measurements is needed to preserve the logical memory for a long time in the presence of measurement and decoherence processes. The threshold for further rounds of syndrome extraction (up to times polynomial in $d$) should be identical. At lower depth, the threshold may shift to higher values as it becomes closer to the code-capacity threshold.
@@
The vertices at the two ends of an error edge in the ${(2+1)D}$ syndrome lattice represent the detection events associated with that simple Pauli error. If that error occurs, the plaquette syndromes of these vertices change sign during the round corresponding to the vertex. Consequently, a vertex that is shared by an odd (even) number of error edges within a subset of all error edges will (will not) exhibit a detection event should this subset of errors occur. The subset of syndrome lattice vertices that are shared by an odd number of error edges within a set of error edges is termed the endpoints of that set. The endpoints of the error edges corresponding to total error, $E$, indicate the detection events linked to these errors. We denote these detection events as $D$.
The decoder’s role is to find a set of errors, denoted as $C$, that would generate the same detection events, $D$, in the syndrome lattice. This set of errors is the proposed correction by the decoder. Our MWPM decoder carries this out by finding a set of error edges with the minimum number of edges that has $D$ as their endpoints. The correction, $C$, is then the set of simple Pauli errors corresponding to these error edges. The combined effect of the occurred errors and the decoder’s correction is linked to the error edges corresponding to $E+C$. Since both the error edges of $E$ and $C$ share $D$ as their endpoints, the error edges corresponding to $E+C$ have no endpoints and form closed paths in the syndrome lattice. Given the periodicity of the syndrome lattice in its hyperbolic $2D$ plane, these ${(2+1)D}$ loops can be either contractible or noncontractible. For each noncontractible loop within the hyperbolic manifold, there exists at least one logical operator that anticommutes with the set of errors of that loop. If $E+C$ comprises an odd number of loops that anticommute with a specific logical operator, the decoding process has resulted in a logical error impacting that logical operator.
@@
Figure 5: Any logical operator error rates for the family of hyperbolic Floquet codes listed in Table 1 under the entangling measurements error model.
@@
Figures 4 and 5 present the MWPM decoder’s performance in correcting the physical errors for the hyperbolic Floquet codes listed in Table 1 11footnotemark: 1, under the phenomenological and entangling measurements error models, respectively. The numerically calculated logical error rates have been determined by running the decoder on the syndrome lattice after applying the physical errors over $3d$ rounds of check measurements, with $d$ the code distance  [^2] Appendix B shows some results where we substitute these $3d$ rounds with $d$ to $4d$ rounds to show that this choice does not affect the results reported here.. In Fig. 4(a) and Fig. 5, an error impacting any of the $2k$ logical operators is considered a failure in a decoding scenario. Conversely, in Fig. 4(b), we focus on solely a specific logical operator and assess decoder’s performance concerning it. Since these codes are based on symmetric graphs (as discussed in Sec. 2), the resulting logical error rate is independent of the specific equivalence class of the chosen logical operator. Hence, Fig. 4(b) would be the same, up to numerical variations, if any other specific logical operator were chosen.
@@
At high physical error rates, both defined logical error rates saturate at a maximum value. This saturation value is 0.5 for the specific logical operator error rate. The reason behind this is that the combined effect of too many errors on a logical operator is equal to an unbiased coin toss: in half of scenarios, there may be an even number of noncontractible loops in $E+C$ that cross this logical operator once, hence cancelling each other’s effects and leaving the logical operator unaffected. For the any logical operator error rate, however, the probability that $E+C$ has no effects on any of the $2k$ logical operators of a code is $1/2^{2k}$, which vanished as $k$ increases. Therefore, this logical error rate saturates at $\sim 1$ for larger codes.
@@
At lower values of the physical error rate, both any logical operator error rate and specific logical operator error rate [panels (a) and (b) in Fig. 4, respectively] show a threshold behavior, as expected from a family of codes with increasing distance. Specifically, for sufficiently low physical error rates, ${p<p_{0}}$, the logical error rate decreases as the code’s physical qubit count increases past $d=4$. The threshold physical error rate, $p_{0}$, for both defined logical error rates is approximately $0.1\,\%$ per measurement round for the phenomenological error model (Fig. 4) and $0.25\,\%$ per measurement round for the entangling measurements error model (Fig. 5). However, it is noticeably smaller for the any logical operator error rate [Fig. 4(a)] compared to the specific logical operator error rate [Fig. 4(b)]. We denote these two values as $p_{0}^{\rm all}$ and $p_{0}^{\rm one}$, respectively, and ${p_{0}^{\rm all}<p_{0}^{\rm one}}$. This subtle difference is a characteristic of finite encoding-rate codes. What links these two defined error rates is the number of logical operators, whereby a code with larger distance and a larger number of physical qubits has more logical operators that could experience a specific logical operator error, resulting in a any logical operator error.
@@
In addition to resulting in different values for the threshold physical error rate, error models also affect the behavior of each code. For example, H64 with a code distance of 4 (defined based on the minimum weight of logical operators during Floquet rounds), behaves similarly to H144 ($d=6$) under the phenomenological error model (Fig. 4), while it only slightly outperforms H16 ($d=2$) under the entangling measurements error model (Fig. 5). This behavior is not unique to the hyperbolic Floquet codes and is observed for the honeycomb Floquet codes as well (see Appendix C). This shows that the true performance of a code under an error model depends on a code distance that is defined for that specific error model.
@@
The threshold physical error rates for hyperbolic codes are typically lower than those for related zero encoding-rate codes [28, 29]. This trend holds true for hyperbolic Floquet codes when compared to honeycomb Floquet codes as well. However, relying solely on the threshold value is not a useful method for comparing different codes. This value pertains to the thermodynamic limit and obscures the actual values of logical error rates, which are more practically relevant. Furthermore, below their threshold rates, hyperbolic codes are storing an extensive number of logical qubits, while the non-hyperbolic codes only preserve a constant number of them. Therefore, the threshold value alone falls short when comparing these codes performing the same task, namely, storing a given number of logical qubits. In response to this, we introduce a new figure of merit that aims at making this comparison.
@@
In Fig. 6 and Fig. 7, we present our proposed comparison between hyperbolic Floquet codes and honeycomb Floquet codes, under the phenomenological and entangling measurements error models, respectively. The four honeycomb Floquet codes featured in these plots, denoted by HC followed by their respective number of physical qubits, are based on symmetric trivalent graphs in Ref. [39] with a girth of six, and are closed hexagonal lattices tessellating the surface of a torus  [^3] Read more about the honeycomb Floquet codes and their performance in Appendix C. Each panel in Fig. 6 and Fig. 7 illustrates the any logical operator error rate of a hyperbolic Floquet code (solid lines) compared to a number of copies of the honeycomb code with the same number of logical qubits (dashed lines). For instance, Fig. 6(a) and Fig. 7(a) compare the H144 code (${[[144,20,6]]}$) with 10 copies of various honeycomb codes. Each of these honeycomb codes stores two logical qubits, and hence, 10 copies have a storage capacity for 20 logical qubits, equal to H144. The total number of physical qubits used in 10 copies of each honeycomb code is noted alongside its respective dashed line. The plotted logical error rate, $p_{L}$, for honeycomb codes is obtained by the reasonable assumption that the logical events in different copies are independent of each other.
@@ 5 Discussion
In this work, we introduced a new family of dynamically generated codes called hyperbolic Floquet codes. These codes are useful for fault-tolerant implementation of multi-qubit quantum memories utilizing weight-two measurements. The hyperbolic nature of these codes enables us to find codes with a finite encoding rate and an increasing distance with the total number of physical qubits. We investigated the fault-tolerance of these codes under two different error models by testing an MWPM decoder on them and showing the existence of a threshold error rate, below which the logical error rates are suppressed. By comparing to honeycomb Floquet codes, we made the case that hyperbolic Floquet codes can achieve logical error suppression better than copies of honeycomb codes storing the same number of logical qubits.
When encoding $k$ logical qubits into $n$ data qubits, we may also use $c$ auxiliary qubits to measure the error syndrome. This means the code uses a total of $n+c$ physical qubits. Thus, the net encoding rate can be defined as ${r=k/(n+c)}$ [12]. For traditional surface code architectures, $c=n-1$, whereas for the newer quantum LDPC codes, $c=n$ [12]. In our scheme, which takes inspiration from the properties of the Floquet code, we have $c=0$ for platforms with native two-body measurements, leading to resource savings. Consequently, our net encoding rate is identical to the regular encoding rate, that is ${k/n=1/8+2/n}$.
@@
In this work, we have focused on the fault-tolerant memory threshold of hyperbolic Floquet codes. For an error correcting code family to be useful for quantum computation it is also necessary to implement state preparation, logical gates, and measurement. State preparation and measurement protocols for our hyperbolic Floquet codes can follow standard protocols from surface codes. Efficiently implementing logical gates with limited time overhead for finite-rate LDPC codes remains a fundamental challenge [20]. However, there has been recent progress adapting lattice surgery methods to LDPC codes that could be applied to our setting [29, 51, 52, 25].
@@ Appendix A Details of Error Detection and Correction
Floquet codes possess the capability to detect potential error occurrences and provide the necessary information for attempting to correct the effects of these errors on logical qubits. Error detection is executed by utilizing the inferred values of the plaquette operators, hereafter referred to as plaquette syndromes or simply syndromes. Each plaquette syndrome is inferred every third round. The inference rounds depend on plaquette’s color; specifically, the green, blue, and red plaquette syndromes are inferred at rounds $r=3n+2$, $33n+$ 3, and $13n+$ 1, respectively, for every nonnegative integer $n$. (No plaquette syndrome is inferred at round ${r=0}$.) By extending a line perpendicular to the hyperbolic plane for each plaquette and marking each syndrome update with a circle on this line, we construct a ${(2+1)D}$ lattice comprising updates of plaquette syndromes, as partially depicted in Fig. 8(a). We refer to this lattice as the syndrome lattice.
In the absence of errors, the plaquette operators become part of the ISG upon their initial inference (rounds $r=2$, $r=3$, and $r=1$ for green, blue, and red plaquettes, respectively) and remain in the ISG, as they commute with all check operators. Consequently, any subsequent inference of a plaquette syndrome yields a value equal to the previously inferred one. This condition is denoted by empty circles in the ${(2+1)D}$ lattice of syndromes in Fig. 8(a). Conversely, an updated plaquette syndrome that differs from its previous value implies the presence of an error (or errors) affecting that syndrome. Following Gidney, we call these events detection events. They are indicated by filled circles in the ${(2+1)D}$ lattice of syndromes.
Figure 8: Detection events for Pauli errors. (a) ${(2+1)D}$ lattice of plaquette syndromes. Each plaquette syndrome is inferred every third measurement round, denoted by empty circles in the timelines of the plaquettes. (b) X and (c) Y Pauli errors between rounds ${r=3n}$ and ${r=3n+1}$ lead to detection events marked on lines perpendicular to the hyperbolic plane. In the presence of errors, inference of a plaquette syndrome may yield a value different than its most recent inferred value. In this case, the inferred syndrome exhibits a detection event, denoted by a filled circle. For the Pauli errors happening between rounds ${r=3n}$ and ${r=3n+1}$, X and Y Paulis are simple Pauli errors and are associated with two detection events, connected to each other by an error edge in the syndrome lattice. A Z Pauli, on the other hand, is a compound Pauli error and is equivalent to a combination of these X and Y Paulis encompassing all four detection events. The four detection events of a compound Pauli error form a hyper-edge in the syndrome lattice.
@@
The lattice’s edge three-coloring ensures that the three measurement checks involving a given physical qubit have distinct Pauli operators associated with that qubit. As a result, a Pauli error on a physical qubit anticommutes with two of those checks. For instance, the X Pauli error on the qubit marked with a circle in Fig. 8(b) anticommutes with the blue (YY) and red (ZZ) checks intersecting at that vertex. Assuming the error occurs between rounds ${r=3n}$ and ${r=3n+1}$, the immediately following blue and red checks will be affected by the error. The blue check’s faulty outcome will be used in the inference of the adjacent red plaquette syndrome resulting in a detection event at round ${r=3n+1}$. This is denoted by a filled circle in Fig. 8(b). At round ${r=3n+2}$, the adjacent green plaquette syndrome will be inferred using two faulty outcomes of both blue and red checks, and hence, will be agnostic to the error. No detection event for the green plaquette syndrome is marked with an empty circle in Fig. 8(b). The blue plaquette syndrome’s inference at round ${r=3n+3}$, however, will use the outcomes from the faulty red check and the unaffected green check, exhibiting a detection event similar to the red plaquette syndrome, indicated with a filled circle. The two detection events of the X Pauli error form an error edge in the syndrome lattice, represented by a dashed line in Fig. 8(b).
@@
The analysis for a Y Pauli error on the same qubit follows similarly. The Y Pauli error on the qubit marked with a circle and its detection events are illustrated in Fig. 8(c). The Y Pauli anticommutes with the green and red checks that intersect at the vertex, while leaving the blue check unaffected. Assuming once again that the error occurs between rounds ${r=3n}$ and ${r=3n+1}$, the blue check’s measurement at round ${r=3n+1}$ is agnostic to the error. However, the red check’s and the green check’s measurements at rounds ${r=3n+2}$ and ${r=3n+3}$ do capture the error. During round ${r=3n+2}$, the green syndrome is updated using the outcomes of the unaffected blue check and the faulty red check, thereby triggering a detection event. In the next round ${r=3n+3}$, the blue syndrome is updated using faulty outcomes from both the red and green checks, leaving it unaffected. Lastly, at round ${r=3n+4}$, the outcomes from the faulty green check and the unaffected blue check lead the red syndrome to exhibit another detection event. These two detection events of the Y Pauli error form an edge in the syndrome lattice, as shown in Fig. 8(c). Notably, for these errors occurring between rounds ${r=3n}$ and ${r=3n+1}$, the two detection events for the Y Pauli error happen with a one round delay as compared to the case of the X Pauli error.
@@
A Z Pauli error that occurs between rounds ${r=3n}$ and ${r=3n+1}$ can be regarded as a combination of X and Y Pauli errors, following from ${Z=-iXY}$. As a result, the associated detection events are also a combination of those corresponding to X and Y Pauli errors. Specifically, a Z Pauli error between rounds ${r=3n}$ and ${r=3n+1}$ leads to the following four detection events: the red syndrome at ${r=3n+1}$, the green syndrome at ${r=3n+2}$, the blue syndrome at ${r=3n+3}$, and once again, the red syndrome at ${r=3n+4}$. These four detection events create a hyper-edge (i.e., an “edge” that connects more than two vertices) in the syndrome lattice.
@@
The decisive factor for a Pauli error to yield either two or four detection events is the last round of measurements. In the outlined example above, the Pauli errors take place between rounds ${r=3n}$ (XX measurements) and ${r=3n+1}$ (YY measurements). In this scenario, X and Y Pauli errors generate two detection events, while the Z Pauli is associated with four. We call Pauli errors with two detection events (X and Y in this case) simple Pauli errors and those with four detection events (Z in this case) compound Pauli errors. During the interval between YY and ZZ measurements, Y and Z Pauli errors are simple, while X Pauli errors are compound. Similarly, between ZZ and XX measurements, Z and X Pauli errors are simple, and Y Pauli errors are compound.

### v4

@@ 3 Error Models and Decoding Method
We employ two error models. A phenomenological model comprising unbiased depolarization errors affecting individual physical qubits and measurement errors associated with the weight-two measurements of the checks, and an entangling measurements model which imitates circuit-level noise for future platforms with native two-qubit measurements, such as Majorana fermions [42]. Considering other noise models and optimizing the code/decoder to the noise is an interesting avenue for future work.
@@
Inferred values of the plaquette operators, called plaquette syndromes or simply syndromes, construct a (2+1)D{(2+1)D} space-time lattice that we refer to as the syndrome lattice. In the absence of errors, consecutive inferences of each plaquette syndrome yield the same value. Conversely, an updated plaquette syndrome that differs from its previous value implies the presence of an error (or errors) affecting that syndrome. We call these events detection events, and they are used in identifying the occurred errors.
Each single-qubit Pauli error, depending on the type of Pauli operator and the most recent round of measurements, yields either two or four detection events. We call Pauli error with two detection events simple Pauli errors and those with four detection events compound Pauli errors. Two-qubit Pauli errors give rise to the detection events corresponding to their constituent single-qubit Paulis, and measurement errors are associated with four detection events. (see Fig. 8 and Fig. 9.)
@@ 4 Fault-Tolerant Memory Threshold
For each hyperbolic Floquet code listed in Table 1, this analysis involves sampling instances of decoding scenarios. In each decoding scenario, depolarization and measurements error channels with strength $p$ (as defined in Sec. 3) are applied during $3d$ rounds of check measurements, corresponding to $d$ Floquet periods, where $d$ is the code distance. Repetition of measurements is needed to preserve the logical memory for a long time in the presence of measurement and decoherence processes. The threshold for further rounds of syndrome extraction (up to times polynomial in $d$) should be identical  [^3] Appendix B shows some results where we substitute these $3d$ rounds with $d$ to $4d$ rounds to show that this choice does not affect the results reported here.. At lower depth, the threshold may shift to higher values as it becomes closer to the code-capacity threshold.
@@
The vertices at the two ends of an error edge in the (2+1)D{(2+1)D} syndrome lattice represent the detection events associated with that simple Pauli error. If that error occurs, the plaquette syndromes of these vertices change sign during the round corresponding to the vertex. Consequently, a vertex that is shared by an odd (even) number of error edges within a subset of all error edges will (will not) exhibit a detection event should this subset of errors occur. The subset of syndrome lattice vertices that are shared by an odd number of error edges within a set of error edges is termed the endpoints of that set. The endpoints of the error edges corresponding to total error, $E$, indicate the detection events linked to these errors. We denote these detection events as $D$.
The decoder’s role is to find a set of errors, denoted as $C$, that would generate the same detection events, $D$, in the syndrome lattice. This set of errors is the proposed correction by the decoder. Our MWPM decoder carries this out by finding a set of error edges with the minimum number of edges that has $D$ as their endpoints. The correction, $C$, is then the set of simple Pauli errors corresponding to these error edges. The combined effect of the occurred errors and the decoder’s correction is linked to the error edges corresponding to $E+C$. Since both the error edges of $E$ and $C$ share $D$ as their endpoints, the error edges corresponding to $E+C$ have no endpoints and form closed paths in the syndrome lattice. Given the periodicity of the syndrome lattice in its hyperbolic $2D$ plane, these (2+1)D{(2+1)D} loops can be either contractible or noncontractible. For each noncontractible loop within the hyperbolic manifold, there exists at least one logical operator that anticommutes with the set of errors of that loop. If $E+C$ comprises an odd number of loops that anticommute with a specific logical operator, the decoding process has resulted in a logical error impacting that logical operator.
@@
Figure 5: Logical error rates for the family of hyperbolic Floquet codes listed in Table 1 under the entangling measurements error model. (a) The any logical operator error rate. (b) The specific logical operator error rate. (See the caption of Fig. 4 for additional information about the two logical error rates.)
@@
Figures 4 and 5 present the MWPM decoder’s performance in correcting the physical errors for the hyperbolic Floquet codes listed in Table 1 22footnotemark: 2, under the phenomenological and entangling measurements error models, respectively. The numerically calculated logical error rates have been determined by running the decoder on the syndrome lattice after applying the physical errors over $3d$ rounds of check measurements, with $d$ the code distance33footnotemark: 3. Then we can examine whether each logical operator has been impacted by an error by counting the number of loops in $E+C$ that anticommute with it. In Fig. 4(a) and Fig. 5(a), an error impacting any of the $2k$ logical operators is considered a failure in a decoding scenario. Conversely, in Fig. 4(b) and Fig. 5(b), we focus on solely a specific logical operator and assess decoder’s performance concerning it. Since these codes are based on symmetric graphs (as discussed in Sec. 2), the resulting logical error rate is independent of the specific equivalence class of the chosen logical operator. Hence, Fig. 4(b) and Fig. 5(b) would be the same, up to numerical variations, if any other specific logical operator were chosen.
@@
At high physical error rates, both defined logical error rates saturate at a maximum value. This saturation value is 0.5 for the specific logical operator error rate. The reason behind this is that the combined effect of too many errors on a logical operator is equal to an unbiased coin toss: in half of scenarios, there may be an even number of noncontractible loops in $E+C$ that cross this logical operator once, hence canceling each other’s effects and leaving the logical operator unaffected. For the any logical operator error rate, however, the probability that $E+C$ has no effects on any of the $2k$ logical operators of a code is $1/2^{2k}$, which vanished as $k$ increases. Therefore, this logical error rate saturates at $\sim 1$ for larger codes.
@@
At lower values of the physical error rate, both any logical operator error rate and specific logical operator error rate [panels (a) and (b), respectively, in both figures] show a threshold behavior, as expected from a family of codes with increasing distance. Specifically, for sufficiently low physical error rates, $p<p_{0}$}, the logical error rate decreases as the code’s physical qubit count increases past $d=4$. The threshold physical error rate, $p_{0}$, for both defined logical error rates is approximately $0.1\,\%$ per measurement round for the phenomenological error model (Fig. 4) and $0.2-0.5\,\%$ per measurement round for the entangling measurements error model (Fig. 5). However, it is noticeably smaller for the any logical operator error rate [Fig. 4(a) and Fig. 5(a)] compared to the specific logical operator error rate [Fig. 4(b) and Fig. 5(b)]. We denote these two values as $p_{0}^{\rm all}$ and $p_{0}^{\rm one}$, respectively, and $p_{0}^{\rm all}<p_{0}^{\rm one}$}. This subtle difference is a characteristic of finite encoding-rate codes. What links these two defined error rates is the number of logical operators, whereby a code with larger distance and a larger number of physical qubits has more logical operators that could experience a specific logical operator error, resulting in a any logical operator error.
@@
In addition to resulting in different values for the threshold physical error rate, error models also affect the behavior of each code. For example, H64 with a code distance of 4 (defined based on the minimum weight of logical operators during Floquet rounds), behaves similarly to H144 ($d=6$) under the phenomenological error model (Fig. 4), while it only slightly outperforms H16 ($d=2$) under the entangling measurements error model (Fig. 5). This behavior is not unique to the hyperbolic Floquet codes and is observed for the honeycomb Floquet codes as well (see Appendix D). This shows that the true performance of a code under an error model depends on a code distance that is defined for that specific error model  [^4] See Appendix C for more details in this regard..
@@
The threshold physical error rates for hyperbolic codes are typically lower than those for related zero encoding-rate codes [28, 29]. This trend holds true for hyperbolic Floquet codes when compared to honeycomb Floquet codes as well. However, relying solely on the threshold value is not a useful method for comparing different codes. This value pertains to the thermodynamic limit and obscures the actual values of logical error rates, which are more practically relevant. Furthermore, below their threshold rates, hyperbolic codes are storing an extensive number of logical qubits, while the non-hyperbolic codes only preserve a constant number of them. Therefore, the threshold value alone falls short when comparing these codes performing the same task, namely, storing a given number of logical qubits. In response to this, we introduce a new figure of merit that aims at making this comparison  [^5] Independently introduced in Refs. [12, 46] as well..
@@
In Fig. 6 and Fig. 7, we present our proposed comparison between hyperbolic Floquet codes and honeycomb Floquet codes, under the phenomenological and entangling measurements error models, respectively. The four honeycomb Floquet codes featured in these plots, denoted by HC followed by their respective number of physical qubits, are based on symmetric trivalent graphs in Ref. [40] with a girth of six, and are closed hexagonal lattices tessellating the surface of a torus  [^6] Read more about the honeycomb Floquet codes and their performance in Appendix D.. Each panel in Fig. 6 and Fig. 7 illustrates the any logical operator error rate of a hyperbolic Floquet code (solid lines) compared to a number of copies of the honeycomb code with the same number of logical qubits (dashed lines). For instance, Fig. 6(a) and Fig. 7(a) compare the H144 code ([[144,20,6]]{[[144,20,6]]}) with 10 copies of various honeycomb codes. Each of these honeycomb codes stores two logical qubits, and hence, 10 copies have a storage capacity for 20 logical qubits, equal to H144. The total number of physical qubits used in 10 copies of each honeycomb code is noted alongside its respective dashed line. The plotted logical error rate, $p_{L}$, for honeycomb codes is obtained by the reasonable assumption that the logical events in different copies are independent of each other.
@@ 5 Discussion
In this work, we introduced a new family of dynamically generated codes called hyperbolic Floquet codes. These codes are useful for fault-tolerant implementation of multi-qubit quantum memories utilizing weight-two measurements. The hyperbolic nature of these codes enables us to find codes with a finite encoding rate and an increasing distance with the total number of physical qubits. We investigated the fault-tolerance of these codes under two different error models by testing an MWPM decoder on them and showing the existence of a threshold error rate, below which the logical error rates are suppressed. By comparing to honeycomb Floquet codes, we made the case that, given the same number of physical qubits, hyperbolic Floquet codes can achieve logical error suppression better than copies of honeycomb codes storing the same number of logical qubits. However, a noteworthy disadvantage for the hyperbolic codes is that the decoding task for a hyperbolic code-block of a given number of logical qubits can be more time-consuming than decoding copies of honeycomb code-blocks encoding the same number of logical qubits. Improving and optimizing decoders for hyperbolic codes, similar to other quantum LDPC codes, could reduce this barrier.
When encoding $k$ logical qubits into $n$ data qubits, we may also use $c$ auxiliary qubits to measure the error syndrome. This means the code uses a total of $n+c$ physical qubits. Thus, the net encoding rate can be defined as r=k/(n+c){r=k/(n+c)} [12]. For traditional surface code architectures, c=n1c=n-1, whereas for the newer quantum LDPC codes, $c=n$ [12]. In our scheme, which takes inspiration from the properties of the Floquet code, we have $c=0$ for platforms with native two-body measurements, leading to resource savings. Consequently, our net encoding rate is identical to the regular encoding rate, that is k/n=1/8+2/n{k/n=1/8+2/n}.
@@
In this work, we have focused on the fault-tolerant memory threshold of hyperbolic Floquet codes. For an error correcting code family to be useful for quantum computation it is also necessary to implement state preparation, logical gates, and measurement. State preparation and measurement protocols for our hyperbolic Floquet codes can follow standard protocols from honeycomb Floquet codes [13, 30, 54]. Efficiently implementing logical gates with limited time overhead for finite-rate LDPC codes remains a fundamental challenge [20]. However, there has been recent progress adapting lattice surgery methods to LDPC codes that could be applied to our setting [29, 55, 56, 25].
@@ Appendix A Details of Error Detection and Correction
Floquet codes possess the capability to detect potential error occurrences and provide the necessary information for attempting to correct the effects of these errors on logical qubits. Error detection is executed by utilizing the inferred values of the plaquette operators, hereafter referred to as plaquette syndromes or simply syndromes. Each plaquette syndrome is inferred every third round. The inference rounds depend on plaquette’s color; specifically, the red, green, and blue plaquette syndromes are inferred at rounds $r=3n+2$, $3n+3$, and $3n+1$, respectively, for every nonnegative integer $n$. (No plaquette syndrome is inferred at round r=0{r=0}.) By extending a line perpendicular to the hyperbolic plane for each plaquette and marking each syndrome update with a circle on this line, we construct a (2+1)D{(2+1)D} lattice comprising updates of plaquette syndromes, as partially depicted in Fig. 8(a). We refer to this lattice as the syndrome lattice.
In the absence of errors, the plaquette operators become part of the ISG upon their initial inference (rounds $r=2$, $r=3$, and $r=1$ for red, green, and blue plaquettes, respectively) and remain in the ISG, as they commute with all check operators. Consequently, any subsequent inference of a plaquette syndrome yields a value equal to the previously inferred one. This condition is denoted by empty circles in the (2+1)D{(2+1)D} lattice of syndromes in Fig. 8(a). Conversely, an updated plaquette syndrome that differs from its previous value implies the presence of an error (or errors) affecting that syndrome. Following Gidney, we call these events detection events. They are indicated by filled circles in the (2+1)D{(2+1)D} lattice of syndromes.
Figure 8: Detection events for Pauli errors. (a) (2+1)D{(2+1)D} lattice of plaquette syndromes. Each plaquette syndrome is inferred every third measurement round, denoted by empty circles in the timelines of the plaquettes. (b) X and (c) Y Pauli errors between rounds r=3n{r=3n} and r=3n+1{r=3n+1} lead to detection events marked on lines perpendicular to the hyperbolic plane. In the presence of errors, inference of a plaquette syndrome may yield a value different than its most recent inferred value. In this case, the inferred syndrome exhibits a detection event, denoted by a filled circle. For the Pauli errors happening between rounds r=3n{r=3n} and r=3n+1{r=3n+1}, X and Y Paulis are simple Pauli errors and are associated with two detection events, connected to each other by an error edge in the syndrome lattice. A Z Pauli, on the other hand, is a compound Pauli error and is equivalent to a combination of these X and Y Paulis encompassing all four detection events. The four detection events of a compound Pauli error form a hyper-edge in the syndrome lattice.
@@
The lattice’s edge three-coloring ensures that the three measurement checks involving a given physical qubit have distinct Pauli operators associated with that qubit. As a result, a Pauli error on a physical qubit anticommutes with two of those checks. For instance, the X Pauli error on the qubit marked with a circle in Fig. 8(b) anticommutes with the green (YY) and blue (ZZ) checks intersecting at that vertex. Assuming the error occurs between rounds r=3n{r=3n} and r=3n+1{r=3n+1}, the immediately following green and blue checks will be affected by the error. The green check’s faulty outcome will be used in the inference of the adjacent blue plaquette syndrome resulting in a detection event at round r=3n+1{r=3n+1}. This is denoted by a filled circle in Fig. 8(b). At round r=3n+2{r=3n+2}, the adjacent red plaquette syndrome will be inferred using two faulty outcomes of both green and blue checks, and hence, will be agnostic to the error. No detection event for the red plaquette syndrome is marked with an empty circle in Fig. 8(b). The green plaquette syndrome’s inference at round r=3n+3{r=3n+3}, however, will use the outcomes from the faulty blue check and the unaffected red check, exhibiting a detection event similar to the blue plaquette syndrome, indicated with a filled circle. The two detection events of the X Pauli error form an error edge in the syndrome lattice, represented by a dashed line in Fig. 8(b).
@@
The analysis for a Y Pauli error on the same qubit follows similarly. The Y Pauli error on the qubit marked with a circle and its detection events are illustrated in Fig. 8(c). The Y Pauli anticommutes with the red and blue checks that intersect at the vertex, while leaving the green check unaffected. Assuming once again that the error occurs between rounds r=3n{r=3n} and r=3n+1{r=3n+1}, the green check’s measurement at round r=3n+1{r=3n+1} is agnostic to the error. However, the blue check’s and the red check’s measurements at rounds r=3n+2{r=3n+2} and r=3n+3{r=3n+3} do capture the error. During round r=3n+2{r=3n+2}, the red syndrome is updated using the outcomes of the unaffected green check and the faulty blue check, thereby triggering a detection event. In the next round r=3n+3{r=3n+3}, the green syndrome is updated using faulty outcomes from both the blue and red checks, leaving it unaffected. Lastly, at round r=3n+4{r=3n+4}, the outcomes from the faulty red check and the unaffected green check lead the blue syndrome to exhibit another detection event. These two detection events of the Y Pauli error form an edge in the syndrome lattice, as shown in Fig. 8(c). Notably, for these errors occurring between rounds r=3n{r=3n} and r=3n+1{r=3n+1}, the two detection events for the Y Pauli error happen with a one round delay as compared to the case of the X Pauli error.
@@
A Z Pauli error that occurs between rounds r=3n{r=3n} and r=3n+1{r=3n+1} can be regarded as a combination of X and Y Pauli errors, following from Z=iXY{Z=-iXY}. As a result, the associated detection events are also a combination of those corresponding to X and Y Pauli errors. Specifically, a Z Pauli error between rounds r=3n{r=3n} and r=3n+1{r=3n+1} leads to the following four detection events: the blue syndrome at r=3n+1{r=3n+1}, the red syndrome at r=3n+2{r=3n+2}, the green syndrome at r=3n+3{r=3n+3}, and once again, the blue syndrome at r=3n+4{r=3n+4}. These four detection events create a hyper-edge (i.e., an “edge” that connects more than two vertices) in the syndrome lattice.
@@
The decisive factor for a Pauli error to yield either two or four detection events is the most recent round of measurements. In the outlined example above, the Pauli errors take place between rounds r=3n{r=3n} (XX measurements) and r=3n+1{r=3n+1} (YY measurements). In this scenario, X and Y Pauli errors generate two detection events, while the Z Pauli is associated with four. We call Pauli errors with two detection events (X and Y in this case) simple Pauli errors and those with four detection events (Z in this case) compound Pauli errors. During the interval between YY and ZZ measurements, Y and Z Pauli errors are simple, while X Pauli errors are compound. Similarly, between ZZ and XX measurements, Z and X Pauli errors are simple, and Y Pauli errors are compound.